The problem is write a equation to show the relationship between n, the number of T-Shirts printed, and C, the total cost in dollars:
T- shirts printing cost
# of T-Shirts Total Cost ($)
20 / 110
30 / 135
40 / 160
50 / 185
60 / 210

Answer by richwmiller(17219)

y=mx+b
135-110=m*(30-20)
25=10m
m=2.5
110=2.5*20+b
b=60
y = 2.5x+60
C = 2.5n+60